# thanks: This work is based on the author’s M.Sc. dissertation (Ref.[1]), but further correcting and expanding upon it

### v1

@@ Abstract
We derive the closed-form one-loop Euler–Heisenberg effective actions for Dirac fermions coupled simultaneously to classical electromagnetic vector and massive pseudo-vector backgrounds within a controlled quasi-static approximation. Through complete diagonalization of the functional Hessian, we systematically delineate the parameter space into distinct sectors characterized by stability properties and spectral structure. We identify subspaces that encompass and extend results from previous studies into a broader class, admitting propagating axial fields as physically viable regimes; strikingly, we note a sector presenting chirality-asymmetric instability. This addresses long-standing questions regarding the well-defined nature, diagonalizability, and stability of the model. From the effective action, we derive novel nonperturbative pair-production rates for simultaneously propagating electromagnetic and axial vector backgrounds; remarkably, we find pronounced vacuum stabilization compared to previous results. Furthermore, we demonstrate that this framework allows for a unified derivation of the chiral anomaly structures in the general case and show that the electromagnetic coupling induces instanton-like configurations for the axial field, even when it is not a fundamental gauge field. As a proof-of-concept, we analyze a cosmological toy model of baryogenesis driven by an axial vector, providing numerical estimates that support the viability of this hypothesis. Additionally, we outline qualitative predictions for Weyl/Dirac semi-metals and briefly discuss potential applications in related phenomena, such as the Strong-CP problem.
@@ I Introduction
The nonperturbative response of the quantum vacuum to strong classical fields is a cornerstone of quantum electrodynamics (QED). Integrating out charged fermions yields the Euler–Heisenberg (EH) effective action [2], which encodes low-energy ($\leq mc^{2}$) nonlinear photon interactions and underpins quantitative predictions for Schwinger pair production [3], vacuum birefringence, photon-photon scattering [4], and related phenomena. For homogeneous electromagnetic backgrounds, closed-form EH expressions are routinely derived via proper-time [5] or numerical worldline [6] methods, establishing the standard framework for field-induced particle production and nonlinear optics [7, *exactplanewave, *exactselfdual1, *exactselfdual2, *exactsauter, *exactplanewave2]. In recent decades, this formalism has been extended to include fermions coupled simultaneously to a vector potential $A_{\mu}$ and a parity-violating axial-vector field $S_{\mu}$—a configuration that arises in three distinct physical arenas: (i) gravitational theories with torsion, where the pseudo-trace of the contorsion tensor acts as $S_{\mu}$ and can seed baryogenesis [13, *stringtorsion, 15, *ShapiroBelyaevTorsion, 17, *KalbRamond, *2Bornot2B]; (ii) high-energy extensions of the Standard Model, where axial fields may catalyze magnetic monopole pair production [20, *monopole]; and (iii) Weyl/Dirac semimetals, where momentum-space node separations modulated by strain or magnetic textures generate emergent axial couplings [22, *torsionweylmetal, *Topologicalmaterials]. Despite their disparate origins, these contexts share a common field-theoretic core, motivating a unified nonperturbative treatment capable of bridging cosmological dynamics and condensed-matter experiments.
@@
Pioneering work by Maroto [25, 26] first revealed that constant axial backgrounds in quasi-static configurations ($\partial_{\mu}S^{\mu}=0$) can suppress Schwinger pair production—a stark departure from pure QED. Subsequent studies generalized the EH action to vector-axial systems, but often yielded operator-valued expressions requiring further contractions, limiting practical utility [27]. Recent advances identified special field configurations—e.g., constant, light-like, or parallel $S_{\mu}$—that admit exact diagonalization and exhibit either enhanced or suppressed pair creation [26, 28, *Copinger1, 30]. Concurrently, condensed-matter theory has clarified how strain and Floquet driving emulate axial fields in topological semimetals, opening pathways to test Lorentz violation and nonlinear axial responses in lab [31, 32, *Diractorsion, 34, *Lorentzviolationsemimetals2]. Yet a systematic classification of the full parameter space—delineating when the model is stable, Hermitian, and diagonalizable—has remained absent. The core obstacle lies in the functional Hessian: axial-vector interference can render it non-Hermitian or defective, signaling vacuum instability. Moreover, consistency conditions from torsion-motivated UV completions (e.g., Planck-scale mass, ghost-freedom [36]) further constrain viable regimes, demanding a rigorous sector-by-sector analysis of physical admissibility.
@@
In this work, we present a complete one-loop analysis of models containing fermions simultaneously coupled to a vector and a possibly massive pseudo-vector fields in configurations with constant field invariants (quasi-static/adiabatic settings), which we shall call electroaxial theories in what follows. Our work builds upon and extends the foundations laid before through the Following contributions: (i) The Complete Hessian diagonalization and systematic parameter-space classification of well-defined and stable subspaces; (ii) The derivation of Exact closed-form effective actions within the viable sectors even for propagating axial-vector field; (iii) The unified derivation of anomaly structures and demonstration of constraints that forces the pseudo-vector to inherit topological structures from the gauge vector field; (iv) The extraction of novel Pair-production rates for propagating axial-vector backgrounds and (v) The Phenomenological applications with numerical estimates for relic field strength from a proof-of-concept toy-model for baryogenesis and qualitative assessments for condensed matter and other possible applications. Remarkably, our results indicate a pronounced vacuum-stabilization when the pseudo-vector is present and estimate bounds for the viability of hypotheses in which baryogenesis is driven by an axial-vector.
@@
Throughout, we employ natural units ($\hbar=c=1$), the Minkowski metric signature (+,,,)(+,-,-,-), and adopt conventions consistent with Refs. [37, *QFTQGdoprof] for spinor algebra. Numerical estimates use $M_{\mathrm{Planck}}=1.22\times 10^{19}\,\mathrm{GeV}$ and Planck 2018 cosmological parameters [39].
@@ II The Electroaxial Theory
along with the free axial-vector action [15, *ShapiroBelyaevTorsion]
@@
For simplicity, we assume that the background fields are smooth and confined to a compact spatial domain, with appropriate boundary conditions, within which the quasi-static approximation holds [42]. This assumption is standard for EH-type effective actions [3, 7] as it enables the application of translation-operator manipulations described in Sec. IV while effectively capturing the physics of slowly varying fields. Generalizations to time-dependent backgrounds extend beyond one-loop exactness and typically require resummation techniques [43] or numerical worldline methods [6]. Under this quasi-static approximation, the extended Dirac operator exhibits elliptic properties and possesses a discrete spectrum, thereby ensuring it is trace-class [5, *HeatKernel2]. Consequently, the traces of $\not${D} and its transpose are equivalent, such that
@@
The transpose of the Dirac operator can be derived using the charge conjugation matrix defined as $C=\gamma^{2}\gamma^{0}$. This matrix exhibits the properties:
@@
In this equation, we denote the quantum corrections as $\overline{\Gamma}\left[A,S\right]$. Direct substitution gives:
@@
Next, we map this result into operator formalism as follows, where we adopt the hat notation for operators temporarily. Let $\hat{x}^{\mu}$ and $\hat{p}_{\mu}$ be the position and momentum operators respectively which obey the commutation relations $\left[\hat{x}^{\mu},\hat{p}_{\nu}\right]=i\delta^{\mu}_{\nu}$, each possessing a complete eigenbasis $\hat{x}^{\mu}\left|x\right>=x^{\mu}\left|x\right$> and $\hat{p}_{\mu}\left|p\right>=p_{\mu}\left|p\right$> satisfying
Next, we transition to an operator formalism, employing hat notation for operators temporarily. Let $\hat{x}^{\mu}$ and $\hat{p}_{\mu}$ be the position and momentum operators, respectively, which satisfy the commutation relations $\left[\hat{x}^{\mu},\hat{p}_{\nu}\right]=i\delta^{\mu}_{\nu}$, each possessing a complete eigenbasis $\hat{x}^{\mu}\left|x\right>=x^{\mu}\left|x\right$> and $\hat{p}_{\mu}\left|p\right>=p_{\mu}\left|p\right$> satisfying
$\left<p\vphantom{x}\right|\left.x\vphantom{p}\right>=\frac{1}{\left(2\pi\right)^{2}}e^{-ip\cdot x}$ (15)
@@
$\left<x\right|\hat{p}_{\mu}\left|\Psi\right>=-i\partial_{\mu}\left<x\vphantom{\Psi}\right|\left.\Psi\vphantom{x}\right$> (16)
@@
where $\hat{\Pi}_{\mu}=\hat{p}_{\mu}-e\hat{A}_{\mu}$, and we will drop the hat notation for the remainder of this paper. By employing the identities [45]:
@@
$\overline{\Gamma}[A,S]=-\frac{i}{2}\mathrm{Tr}\ln\left[\frac{\left(\Pi_{\mu}+\eta S_{\mu}\gamma^{5}\right)^{2}-\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}-m^{2}}{p^{2}-m^{2}}\right],$ (19)
@@
$\overline{\Gamma}\left[A,S\right]=\frac{i}{2}\int_{0}^{\infty}\frac{d\tau}{\tau}e^{-i\tau m^{2}}\mathrm{Tr}\left[e^{-i\tau\mathcal{H}}-e^{-i\tau p^{2}}\right]$ (20)
@@
$\mathcal{H}=\left(\Pi_{\mu}+\eta S_{\mu}\gamma^{5}\right)^{2}-\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}$ (21)
This representation brings forth a long and well discussed [3, 41, 38] analogy between these exponentials in Eq. (20) and time-evolution operators such as $e^{-i\tau\mathcal{H}}\left|x\right>=\left|x;\tau\right$>. In this context, $\tau$ acts as a form of “proper time,” while the remaining terms function akin to a “Hamiltonian” that adheres to its own Schrödinger-like functional equations, expressed as follows:
@@
$\mathrm{Tr}\ e^{-i\tau p^{2}}=\int d^{4}x\left<x\right|e^{-i\tau p^{2}}\left|x\right>\mathrm{tr}\mathbf{1}=\lim_{y\to x}\frac{4}{(2\pi)^{4}}\int d^{4}x\ d^{4}p\left<y;0\vphantom{p}\right|\left.p\vphantom{y;0}\right>\left<p\vphantom{x;\tau}\right|\left.x;\tau\vphantom{p}\right>=\int d^{4}x\ \frac{i}{(2\pi\tau)^{2}}$ (23)
@@ III Hessian Diagonalization and Sector Classification
We can recast (21) as $\mathcal{H}=\left(\Pi^{2}+\eta^{2}S^{2}\right)\mathbf{1}+\mathcal{H}_{\mathrm{R}}$. The remaining term is defined as:
$\mathcal{H}_{\mathrm{R}}=2\eta\Pi_{\mu}S^{\mu}\gamma^{5}-\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}$ (26)
@@
From Eqs. (28) and (103), we observe that there exist certain points in the parameter space at which the Hessian ceases to be Hermitian. Specifically, this occurs when $\mathcal{D}_{-}\to 0$ or $\infty$, indicating configurations where (i) $\Delta\to 0$ [Eq. 104], (ii) $\Delta\to\infty$, (iii) $\mathcal{A}\to\infty$, and (iv) $\mathcal{B}\to\infty$ are unstable, suggesting potential symmetry breaking or phase transitions [49, *degeneracies, *PhaseTransitions]. Additionally, there are regions where the imaginary part of the eigenvalues may be relevant or even dominant. While this may initially seem unphysical for effective actions derived from fundamental theories (such as Quantum Gravity and Extensions of the Standard Model), it could prove beneficial for deriving particle creation rates within these theories, or be applicable in other domains, such as Open Systems [52] and PT-Symmetric models [53], which are useful in Quantum Optics [54], Nuclear Physics [55, *ComplexHamiltonianApplication], and beyond.
@@
In this regime, the conditions $\Pi_{\mu}S^{\mu}=\widetilde{V}^{\mu\nu}S_{\nu}=0$ and $S^{2}\simeq 0$ hold, leading to the simplification of the Hessian to $\mathcal{H}_{\mathrm{R}}=-\frac{1}{2}V_{\mu\nu}\sigma^{\mu\nu}$. This yields expressions that are analogous to those found in QED, where the quantity $V^{\mu\nu}$ serves as an extended version of the Faraday tensor $F_{\mu\nu}$. The eigenvalues associated with this system can be expressed as follows:
@@
In the MT case, the fields are constrained by the condition $\Pi_{\mu}S^{\mu}=0$, indicating that they are transverse and orthogonal, specifically $A_{\mu}S^{\mu}=\partial_{\mu}\left(S^{\mu}\right)=0$. Under these constraints, the eigenvalues take the following form:
@@
The system remains well-behaved as long as the condition $\mathcal{V}\geq 2m^{2}\eta^{2}S^{2}\pm\Re\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}$ is satisfied. It is noteworthy that the EQED-type is encompassed within this broader subspace.
@@
$\frac{\delta\mathrm{S}\left[A,S\right]}{\delta S_{\mu}}\approx S^{\mu}+\frac{1}{M^{2}_{S}}j_{5}^{\mu}=0$ (33)
where $j_{5}^{\mu}=\left<\,\overline{\!{\psi}}\gamma^{\mu}\gamma^{5}\psi\right$>. This leads us to conclude that:
$\left|\Pi_{\mu}S^{\mu}\right|\sim\frac{m^{2}}{M^{2}_{S}}\approx 0$ (34)
From this, we can further simplify the eigenvalues to:
@@
These eigenvalues remain stable under the condition $\mathcal{V}\geq 2\eta m\left(\eta mS^{2}+\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\right)$ and are amenable to analysis since they avoid the complication of handling $\left(\Pi_{\mu}S^{\mu}\right)$ within a square root.
@@
This subspace is contained within the previous one, but it introduces the additional constraint of $\mathcal{V}=0$. In a teleparallel or Einstein–Cartan setting, where the axial vector is derived from a torsion tensor ($S_{\mu}=\varepsilon_{\mu\nu\alpha\beta}T^{\mu\nu\alpha\beta}$ [15, *ShapiroBelyaevTorsion]), this constraint realizes a Rainich-style [58] identification of an electromagnetic-like structure with torsion. Notably, this constraint also necessitates the self-duality of the fields, expressed as $\widetilde{V}^{\mu\nu}=\pm iV^{\mu\nu}$.
@@
When the condition $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|=0$ is satisfied, the Hessian of a system lying within this subspace can be hermitian if and only if $mS\leq\left|\Pi_{\mu}S^{\mu}\right|$. A special simplification arises for a pseudo-vector in light-like configurations ($S^{2}=0$), which can be achieved either (i) classically if it is massless or (ii) off-shell if it is quantized. In such cases, the Hessian takes the form:

### v2

@@ Abstract
We derive the closed-form one-loop Euler–Heisenberg effective actions for Dirac fermions coupled simultaneously to classical electromagnetic vector and massive pseudo-vector backgrounds within a controlled quasi-static approximation. Through complete diagonalization of the functional Hessian, we systematically delineate the parameter space into distinct sectors characterized by stability properties and spectral structure. We identify subspaces that encompass and extend results from previous studies into a broader class, admitting propagating axial fields as physically viable regimes; strikingly, we note sectors presenting chirality-asymmetric instability. This addresses long-standing questions regarding the well-defined nature, diagonalizability, and stability of these models. From the effective actions, we derive novel nonperturbative pair-production rates for simultaneously propagating electromagnetic and axial vector backgrounds; remarkably, we find pronounced vacuum stabilization compared to previous results. Furthermore, we demonstrate that this framework allows for a unified derivation of the chiral anomaly structures in the general case and show that the electromagnetic coupling induces instanton-like configurations for the axial field, even when it is not a fundamental gauge field. As a proof-of-concept, we analyze a cosmological toy model of baryogenesis driven by an axial vector, providing numerical estimates that support the viability of this hypothesis. Additionally, we outline qualitative predictions for Weyl/Dirac semi-metals and briefly discuss potential applications in related phenomena, such as the Strong-CP problem.
@@ I Introduction
The nonperturbative response of the quantum vacuum to strong classical fields is a cornerstone of quantum electrodynamics (QED). Integrating out charged fermions yields the Euler–Heisenberg (EH) effective action [2], which encodes low-energy ($\leq mc^{2}$) nonlinear photon interactions and underpins quantitative predictions for Schwinger pair production [3], vacuum birefringence, photon-photon scattering [4], and related phenomena. For homogeneous electromagnetic and/or scalar backgrounds, closed-form EH expressions are routinely derived via resummation [5, *HeatKernel2, 7, *resummedeffectiveaction, *sebastianresummedheatkerneleffective] or worldline [10] methods, establishing the standard framework for field-induced particle production and nonlinear optics [11, *exactplanewave, *exactselfdual1, *exactselfdual2, *exactsauter, *exactplanewave2, *nonlocalEffActionSebastian]. In recent decades, this formalism has been extended to include fermions coupled simultaneously to a vector potential $A_{\mu}$ and a parity-violating axial-vector field $S_{\mu}$—a configuration that arises in three distinct physical arenas: (i) gravitational theories with torsion, where the pseudo-trace of the contorsion tensor acts as $S_{\mu}$ and can seed baryogenesis [18, *stringtorsion, 20, *ShapiroBelyaevTorsion, 22, *KalbRamond, *2Bornot2B]; (ii) high-energy extensions of the Standard Model, where axial fields may catalyze magnetic monopole pair production [25, *monopole]; and (iii) Weyl/Dirac semimetals, where momentum-space node separations modulated by strain or magnetic textures generate emergent axial couplings [27, *torsionweylmetal, *Topologicalmaterials]. Despite their disparate origins, these contexts share a common field-theoretic core, motivating a unified nonperturbative treatment capable of bridging cosmological dynamics and condensed-matter experiments.
@@
Pioneering work by Maroto [30, 31] first revealed that constant axial backgrounds in quasi-static configurations ($\partial_{\mu}S^{\mu}=0$) can suppress Schwinger pair production—a stark departure from pure QED. Subsequent studies generalized the EH action to vector-axial systems [32, *LorentzbreakingEHPetrov, *bastianelliunifiedworldlineaxial, *LorentzbreakingEHtensorPetrov], but often yielded operator-valued expressions requiring further contractions, limiting practical utility. Recent advances identified special field configurations—e.g., constant, light-like, or parallel $S_{\mu}$—that admit exact diagonalization and exhibit either enhanced or suppressed pair creation [31, 36, *Copinger1, *Sebastianheatkernelsresummationsspinor, 39]. Concurrently, condensed-matter theory has clarified how strain and Floquet driving emulate axial fields in topological semimetals, opening pathways to test Lorentz violation and nonlinear axial responses in lab [40, 41, *Diractorsion, 43, *Lorentzviolationsemimetals2]. Yet a systematic classification of the full parameter space—delineating when the model is stable, Hermitian, and diagonalizable—has remained absent. The core obstacle lies in the functional Hessian: axial-vector interference can render it non-Hermitian or defective, signaling vacuum instability. Moreover, consistency conditions from torsion-motivated UV completions (e.g., Planck-scale mass, ghost-freedom [45]) further constrain viable regimes, demanding a rigorous sector-by-sector analysis of physical admissibility.
@@
In this work, we present a complete one-loop analysis of models containing fermions simultaneously coupled to a vector and a possibly massive pseudo-vector fields in configurations with constant field invariants (quasi-static/adiabatic settings), which we shall call electroaxial theories in what follows. Our work builds upon and extends the foundations laid before through the Following contributions: (i) The Complete Hessian diagonalization and systematic parameter-space classification of well-defined and stable subspaces; (ii) The derivation of Exact closed-form effective actions within the viable sectors even for propagating axial-vector field; (iii) The unified derivation of anomaly structures and demonstration of constraints that forces the pseudo-vector to inherit topological structures from the gauge vector field; (iv) The extraction of novel Pair-production rates for propagating axial-vector backgrounds and (v) The Phenomenological applications with numerical estimates for relic field strength from a proof-of-concept toy-model for baryogenesis and qualitative assessments for condensed matter and other possible applications. Remarkably, our results indicate the possibility of chirality-asymmetric instability; a pronounced vacuum-stabilization when the pseudo-vector is present and estimate bounds for the viability of hypotheses in which baryogenesis is driven by an axial-vector.
@@
Throughout, we employ natural units ($\hbar=c=1$), the Minkowski metric signature (+,,,)(+,-,-,-), and adopt conventions consistent with Refs. [46, 47, *QFTQGdoprof] for spinor algebra. Numerical estimates use $M_{\mathrm{Planck}}=1.22\times 10^{19}\,\mathrm{GeV}$ and Planck 2018 cosmological parameters [49].
@@ II The Electroaxial Theory
along with the free axial-vector action [20, *ShapiroBelyaevTorsion]
@@
For simplicity, we assume that the background fields are smooth and confined to a compact spatial domain, with appropriate boundary conditions, within which the quasi-static approximation holds [51]. This assumption is standard for EH-type effective actions [3, 11] as it enables the application of translation-operator manipulations described in Sec. IV while effectively capturing the physics of slowly varying fields. Generalizations to time-dependent backgrounds extend beyond one-loop exactness and typically require resummation techniques [5, *HeatKernel2, 7, *resummedeffectiveaction, *sebastianresummedheatkerneleffective] or worldline methods [10]. Under this quasi-static approximation, the extended Dirac operator exhibits elliptic properties and possesses a discrete spectrum, thereby ensuring it is trace-class [5, *HeatKernel2]. Consequently, the traces of $\not${D} and its transpose are equivalent, such that
@@
The transpose of the Dirac operator can be derived using the charge conjugation matrix defined as $C=i\gamma^{2}\gamma^{0}$. This matrix exhibits the properties:
@@
In this equation, we denote the quantum corrections as $\overline{\Gamma}[A,S]$. Direct substitution gives:
@@
Next, we transition to an operator formalism, employing hat notation for operators temporarily. Let $\hat{x}^{\mu}$ and $\hat{p}_{\mu}$ be the position and momentum operators, respectively, which satisfy the commutation relations $\hat{x}^{\mu},\hat{p}_{\nu}]=-i\delta^{\mu}_{\nu}$, each possessing a complete eigenbasis $\hat{x}^{\mu}\left|x\right>=x^{\mu}\left|x\right$> and $\hat{p}_{\mu}\left|p\right>=p_{\mu}\left|p\right$> satisfying
$\left<p\vphantom{x}\right|\left.x\vphantom{p}\right>=\frac{1}{\left(2\pi\right)^{2}}e^{ip\cdot x}$ (15)
@@
$\left<x\right|\hat{p}_{\mu}\left|\Psi\right>=i\partial_{\mu}\left<x\vphantom{\Psi}\right|\left.\Psi\vphantom{x}\right$> (16)
@@
where $\hat{\Pi}_{\mu}=\hat{p}_{\mu}+e\hat{A}_{\mu}$, and we will drop the hat notation for the remainder of this paper. By employing the identities [52]:
@@
$\overline{\Gamma}[A,S]=-\frac{i}{2}\mathrm{Tr}\ln\left[\frac{\left(\Pi_{\mu}+\eta S_{\mu}\gamma^{5}\right)^{2}+\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}-m^{2}}{p^{2}-m^{2}}\right]$ (19)
@@
$\overline{\Gamma}\left[A,S\right]=\frac{i}{2}\int_{0}^{\infty}\frac{d\tau}{\tau}e^{-i\tau m^{2}}\mathrm{Tr}\left[e^{i\tau\mathcal{H}}-e^{i\tau p^{2}}\right]$ (20)
@@
$\mathcal{H}=\left(\Pi_{\mu}+\eta S_{\mu}\gamma^{5}\right)^{2}+\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}$ (21)
This representation brings forth a long and well discussed [3, 47, *QFTQGdoprof] analogy between these exponentials in Eq. (20) and time-evolution operators such as $e^{i\tau\mathcal{H}}\left|x\right>=\left|x;\tau\right$>. In this context, $\tau$ acts as a form of “proper time,” while the remaining terms function akin to a “Hamiltonian” that adheres to its own Schrödinger-like functional equations, expressed as follows:
@@
$\mathrm{Tr}\ e^{i\tau p^{2}}=\int d^{4}x\left<x\right|e^{i\tau p^{2}}\left|x\right>\mathrm{tr}\mathbf{1}=\lim_{y\to x}\frac{4}{(2\pi)^{4}}\int d^{4}x\ d^{4}p\left<y;0\vphantom{p}\right|\left.p\vphantom{y;0}\right>\left<p\vphantom{x;\tau}\right|\left.x;\tau\vphantom{p}\right>=\int d^{4}x\ \frac{i}{(2\pi\tau)^{2}}$ (23)
@@ III Hessian Diagonalization and Sector Classification
We can recast (21) as $\mathcal{H}=(\Pi^{2}+\eta^{2}S^{2})\mathbf{1}+\mathcal{H}_{\mathrm{R}}$. The remaining term is defined as:
$\mathcal{H}_{\mathrm{R}}=2\eta\Pi_{\mu}S^{\mu}\gamma^{5}+\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}$ (26)
@@
From Eqs. (28) and (106), we observe that there exist certain points in the parameter space at which the Hessian ceases to be Hermitian. Specifically, this occurs when $\mathcal{D}_{-}\to 0$ or $\infty$, indicating configurations where (i) $\Delta\to 0$ [Eq. 107], (ii) $\Delta\to\infty$, (iii) $\mathcal{A}\to\infty$, and (iv) $\mathcal{B}\to\infty$ are unstable, suggesting potential symmetry breaking or phase transitions [56, *degeneracies, *PhaseTransitions]. Additionally, there are regions where the imaginary part of the eigenvalues may be relevant or even dominant. While this may initially seem unphysical for effective actions derived from fundamental theories (such as Quantum Gravity and Extensions of the Standard Model), it could prove beneficial for deriving particle creation rates within these theories, or be applicable in other domains, such as Open Systems [59] and PT-Symmetric models [60], which are useful in Quantum Optics [61], Nuclear Physics [62, *ComplexHamiltonianApplication], and beyond.
@@
In this regime, the conditions $\Pi_{\mu}S^{\mu}=\widetilde{V}^{\mu\nu}S_{\nu}=0$ and $S^{2}\simeq 0$ hold, leading to the simplification of the Hessian to $\mathcal{H}_{\mathrm{R}}=\frac{1}{2}V_{\mu\nu}\sigma^{\mu\nu}$. This yields expressions that are analogous to those found in QED, where the quantity $V^{\mu\nu}$ serves as an extended version of the Faraday tensor $F_{\mu\nu}$. The eigenvalues associated with this system can be expressed as follows:
@@
In the MT case, the fields are constrained by the condition $\Pi_{\mu}S^{\mu}=0$, indicating that they are transverse and orthogonal, specifically $A_{\mu}S^{\mu}=\partial_{\mu}(S^{\mu})=0$. Under these constraints, the eigenvalues take the following form:
@@
The system remains well-behaved as long as the condition $\mathcal{V}+\Re\sqrt{4m^{2}\eta^{2}(\widetilde{V}^{\mu\nu}S_{\nu})^{2}-\mathcal{W}^{2}}\geq 2m^{2}\eta^{2}S^{2}$ is satisfied. Three special cases are noteworthy within this subspace: First, the EQED-type is encompassed within this broader subspace. Second, If $\mathcal{V}=\mathcal{W}=0$, Eq. (31) implies
$\lambda^{(\pm)}_{\pm}=\left(\pm\right)2\sqrt{\eta m\left(-\eta mS^{2}\pm\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\right)}$ (32)
When the pseudo-vector is light-like ($S^{2}=0$), the full Hessian becomes:
$\mathcal{H}=\begin{bmatrix}\Pi^{2}\mathbf{1}+2\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\,\sigma_{z}&0\\ 0&\Pi^{2}\mathbf{1}+2i\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\,\sigma_{z}\end{bmatrix}$ (33)
where $\sigma_{z}=\mathrm{diag}\{1,-1\}$ represent the Pauli matrix in the $z$-direction. Third, when $(\widetilde{V}^{\mu\nu}S_{\nu})^{2}=0$ and the generalized fields are (anti-)self-dual, i.e., $\mathcal{V}=\pm i\mathcal{W}$, it follows that
$\lambda^{(\pm)}_{\pm}=(\pm)\sqrt{2\left(1\pm 1\right)\mathcal{V}-4m^{2}\eta^{2}S^{2}}$ (34)
which means the Hessian can show a similar structure to Eq. (33) provided $\mathcal{V}\geq m^{2}\eta^{2}S^{2}$. It can be seem that Eqs. (32), (33) and (34) manifest a chiral-asymmetric instability, offering compelling realizations of Very Special Relativity (VSR). This framework will be examined in greater detail in subsequent paragraphs of this section.
@@
$\frac{\delta\mathrm{S}\left[A,S\right]}{\delta S_{\mu}}\approx S^{\mu}+\frac{\eta}{M^{2}_{S}}j_{5}^{\mu}=0$ (36)
where $j_{5}^{\mu}=\langle\,\overline{\!{\psi}}\gamma^{\mu}\gamma^{5}\psi\rangle$. This leads us to conclude that:
$\left|\Pi_{\mu}S^{\mu}\right|\propto\frac{m}{M^{2}_{S}}\approx 0$ (37)
From this relation, we can further simplify the eigenvalues, yielding:
@@
If we take into account the specific condition under which $|V^{\mu\nu}S_{\nu}|=0$, we can simplify this expression even further to:
$\lambda^{(\pm)}_{\pm}\approx\left(\pm\right)\left(\sqrt{2\mathcal{V}}\pm 2i\eta mS\right)$ (39)
This reduction is justified by the relation that connects the quantities involved, specifically $(\widetilde{V}^{\mu\nu}S_{\nu})^{2}=-2S^{2}\mathcal{V}+(V^{\mu\nu}S_{\nu})^{2}$. These eigenvalues retain some degree of stability under the condition $\mathcal{V}+2\eta m|\widetilde{V}^{\mu\nu}S_{\nu}|\geq 2\eta^{2}m^{2}S^{2}$ and are amenable to analysis since they avoid the complication of handling $(\Pi_{\mu}S^{\mu})$ within a square root. In addition, under the condition that $\Pi_{\mu}S^{\mu}$ is not negligible while $|V^{\mu\nu}S_{\nu}|=0$, we obtain the relation
$\lambda^{(\pm)}_{\pm}=\left(\pm\right)\left[2\eta\sqrt{\left(\Pi_{\mu}S^{\mu}\right)^{2}-m^{2}S^{2}}\pm\sqrt{2\mathcal{V}}\right]$ (40)
This expression can be further simplified in the case where the pseudovector takes on a light-like configuration.
@@
This subspace is contained within the previous one, but it introduces the additional constraint of $\mathcal{V}=0$. In a teleparallel or Einstein–Cartan setting, where the axial vector is derived from a torsion tensor [20, *ShapiroBelyaevTorsion] ($S_{\mu}=\varepsilon_{\mu\nu\alpha\beta}T^{\mu\nu\alpha\beta}$), this constraint realizes a Rainich-style [65] identification of an electromagnetic-like structure with torsion. Notably, this constraint also necessitates the self-duality of the fields, expressed as $\widetilde{V}^{\mu\nu}=\pm iV^{\mu\nu}$.
@@
When the condition $|\widetilde{V}^{\mu\nu}S_{\nu}|=0$ is satisfied, the Hessian of a system lying within this subspace can be hermitian if and only if $mS\leq|\Pi_{\mu}S^{\mu}|$. A special simplification arises for a pseudo-vector in light-like configurations ($S^{2}=0$), which can be achieved either (i) classically if it is massless or (ii) off-shell if it is quantized. In such cases, the Hessian takes the form: